Write the following sets in the roaster form.
 
(i) A = {x | x is a positive integer less than 10 and 2x – 1 is an odd number}
 
(ii) C = {x : x2 + 7x – 8 = 0, x ∈ R}

Grade:12

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

(i) 2^x– 1 is always an odd number for all positive integral values of x since 2xis an even number.

In particular, 2x– 1 is an odd number for x = 1, 2, … , 9.

Therefore, A = {1, 2, 3, 4, 5, 6, 7, 8, 9}

(ii) x^2+ 7x – 8 = 0

(x + 8) (x – 1) = 0

x = – 8 or x = 1

Therefore, C = {– 8, 1}

Thanks